# Beware of the Classical Benchmark Instances for the Traveling Salesman Problem with Time Windows

### v1

@@ 4.2 Results for the TSPTW-D
For the TSPTW-D, we set a time limit of $30$ minutes per instance. Table 3 summarizes the results for the benchmarks Asc, DaS, Gen, Ohl, and Ohl-D. It includes the results of Algorithm 3 applied to the transport network $D$ (columns labeled Algorithm 3) and its reverse network $D^{-1}$ (columns labeled Algorithm 3-1), as the method is not symmetric, and the running times differ considerably between the two. For comparison, Table 3 also reports results from TilkIrnich2017 and Lera-RomeroMirandaBrontSoulignac2022.
We do not consider the benchmarks Pot and Pes because their travel times are rounded to $10^{-4}$, and Algorithm 3 cannot handle the large number of TSPTW-M instances in the range $10^{4}T]$. This also applies to the four unsolved instances in Asc (rbg021.xx for $x\in[\![6,9]\!]$), which were derived from rbg021 by increasing both the horizon and the deadline times in the time windows. Additionally, we exclude the benchmark Lan because its instances are trivial and were not tested by TilkIrnich2017 or Lera-RomeroMirandaBrontSoulignac2022. We also note that TilkIrnich2017; Lera-RomeroMirandaBrontSoulignac2022 did not evaluate their algorithms on the Gen instances with $n\geq 80$ and $\omega\in\{80,100\}$. It is worth mentioning that Lera-RomeroMirandaBrontSoulignac2022 solved all the instances in Pot, while TilkIrnich2017 solved all but three. However, both methods were tested with a time limit of 3 hours per instance.
@@
Table 3 reinforces the conclusions from Section 4: Algorithm 3 (on either $D$ or $D^{-1}$) solves all but one of the classical instances with $50$ or more customers, including the 14 instances that were not solved by TilkIrnich2017 or Lera-RomeroMirandaBrontSoulignac2022. Furthermore, Algorithm 3 outperforms both methods on the larger instances with higher values of $n$.
By design, one might expect Algorithm 3 to be much less efficient than Algorithm 2, as it solves $\Theta(T)$ instances of the TSPTW-M in the worst case. However, as implied by Table 3, only a small number of these instances are actually passed to Algorithm 2 for solving in the benchmark instances. Most feasible departure times $p$ in Loop 611 are either solved by the local search heuristic or ignored because of the best-known solution $R^{*}$. Moreover, $\operatorname{last}_{p}$ is typically small in the largest instances, which helps reduce the number of instances solved by Algorithm 2.
The main conclusion is that the TSPTW-D should be evaluated on “harder” instances, especially those with wider time windows, to better assess the performance of algorithms in more challenging scenarios.
@@ 5 Conclusions
We presented a simple informed search method for the TSPTW-M, which excels in solving all classical benchmark instances with 50 or more customers in under 10 seconds each. However, it struggles to solve even the smallest instances when their time windows are sufficiently loose. By running the algorithm as a preprocessing step for a limited amount of time, any method can solve these instances, meaning they should no longer be used for benchmarking purposes.
An important question that arises from this work is how valid and accurate are the conclusions drawn from methods that have been tested exclusively on these classical benchmark instances? Are these methods truly efficient in a general setting, or are they, perhaps unconsciously, exploiting some inherent bias in these benchmarks? Since these instances can be "hacked" by simpler methods, it is crucial to examine how generalizable the reported results are. Exploring this issue further could help ensure that conclusions drawn from these benchmarks are not only valid for the specific instances tested but also robust across a wider range of real-world problem scenarios.
Our method performs particularly well on the benchmarks Lan, Dum, Gen, Ohl, and DaS, all of which share the common strategy of defining a time window $x-\omega,x+\omega]$ for a vertex visited at time $x$ in a given route. Here, $\omega$ is relatively small when compared to $n$ and $T$. Due to the simplicity of this approach, it is commonly used to generate training datasets for many machine learning methods, which can lead to overfitting and biased conclusions. As noted by RifkiSolnon2025, a simple solution to make these instances more challenging for an informed search algorithm like ours is to extend the time windows. The varying levels of difficulty among these generated instances provide a better foundation for drawing deeper conclusions.
Given the exceptional performance of our method on the larger benchmark instances, it is not surprising that a trivial procedure of repeatedly running the solver across the entire horizon is sufficient to solve the TSPTW-D on most of the instances. The issue is that many of these instances are easy even for the TSPTW-D. Therefore, these instances should also be avoided when evaluating methods for the TSPTW-D. What remains to be explored is whether a similar informed search can be applied to solve the TSPTW-TT on these instances. Nonetheless, we believe it would be wise to evaluate and train future methods on instances with varying levels of time window tightness, even for the TSPTW-TT.

### v2

@@ 4.2 Results for the TSPTW-D
For the TSPTW-D, we set a time limit of $30$ minutes per instance. Table 3 summarizes the results for the benchmarks Asc, DaS, Gen, Ohl, and Ohl-D. It includes the results of Algorithm 3 applied to the transport network $D$ (columns labeled Algorithm 3) and its reverse network $D^{-1}$ (columns labeled Algorithm 3-1), as the method is not symmetric, and the running times differ considerably between the two. For comparison, Table 3 also reports results from Tilk and Irnich (2017) and Lera-Romero et al. (2022).
We do not consider the benchmarks Pot and Pes because their travel times are rounded to $10^{-4}$, and Algorithm 3 cannot handle the large number of TSPTW-M instances in the range $10^{4}T]$. This also applies to the four unsolved instances in Asc (rbg021.xx for $x\in[\![6,9]\!]$), which were derived from rbg021 by increasing both the horizon and the deadline times in the time windows. Additionally, we exclude the benchmark Lan because its instances are trivial and were not tested by Tilk and Irnich (2017) or Lera-Romero et al. (2022). We also note that Tilk and Irnich (2017); Lera-Romero et al. (2022) did not evaluate their algorithms on the Gen instances with $n\geq 80$ and $\omega\in\{80,100\}$. It is worth mentioning that Lera-Romero et al. (2022) solved all the instances in Pot, while Tilk and Irnich (2017) solved all but three. However, both methods were tested with a time limit of 3 hours per instance.
@@
Table 3 reinforces the conclusions from Section 4: Algorithm 3 (on either $D$ or $D^{-1}$) solves all but one of the classical instances with $50$ or more customers, including the 14 instances that were not solved by Tilk and Irnich (2017) or Lera-Romero et al. (2022). Furthermore, Algorithm 3 outperforms both methods on the larger instances with higher values of $n$.
By design, one might expect Algorithm 3 to be much less efficient than Algorithm 2, as it solves $\Theta(T)$ instances of the TSPTW-M in the worst case. However, as implied by Table 3, only a small number of these instances are actually passed to Algorithm 2 for solving in the benchmark instances. Most feasible departure times $p$ in Loop 712 are either solved by the local search heuristic or ignored because of the best-known solution $R^{*}$. Moreover, $\operatorname{last}_{p}$ is typically small in the largest instances, which helps reduce the number of instances solved by Algorithm 2. In short, solving the TSPTW-D on these instances is not substantially harder than solving the TSPTW-M.
The main conclusion is that the evaluation of the TSPTW-D should be complemented with more difficult instances, especially those with wider time windows, to better assess the performance of algorithms in challenging scenarios.
@@ 5 Conclusions
We presented a simple informed search method for the TSPTW-M, which excels in solving all classical benchmark instances with 50 or more customers in under 10 seconds each. However, it struggles to solve even the smallest instances when their time windows are sufficiently loose. By running the algorithm as a preprocessing step for a limited amount of time, any method can solve these instances, meaning it is generally advisable to complement them with harder instances featuring looser time windows, rather than using them alone for benchmarking.
An important question that arises from this work is how valid and accurate are the conclusions drawn from methods that have been tested exclusively on these classical benchmark instances? Are these methods truly efficient in a general setting, or are they, perhaps unconsciously, exploiting some inherent bias in these benchmarks? Since these instances appear easier for simple enumerative methods, it is crucial to examine how generalizable the reported results are. Exploring this issue further could help ensure that conclusions drawn from these benchmarks are not only valid for the specific instances tested but also robust across a wider range of real-world problem scenarios.
Our method performs particularly well on the benchmarks Lan, Dum, Gen, Ohl, and DaS, all of which share the common strategy of defining a time window $x-\omega,x+\omega]$ for a vertex visited at time $x$ in a given route. Here, $\omega$ is relatively small when compared to $n$ and $T$. Due to the simplicity of this approach, it is commonly used to generate training datasets for many machine learning methods, which can lead to overfitting and biased conclusions. As noted by Rifki and Solnon (2025), a simple solution to make these instances more challenging for an informed search algorithm like ours is to extend the time windows. The varying levels of difficulty among these generated instances provide a better foundation for drawing deeper conclusions.
Given the exceptional performance of our method on the larger benchmark instances, it is not surprising that a trivial procedure of repeatedly running the solver across the entire horizon is sufficient to solve the TSPTW-D on most of the instances. The issue is that many of these instances are easy even for the TSPTW-D. Therefore, these instances are not sufficient when evaluating methods for the TSPTW-D. What remains to be explored is whether a similar informed search can be applied to solve the TSPTW-TT on these instances. Nonetheless, we believe it would be wise to evaluate and train future methods on instances with varying levels of time window tightness, even for the TSPTW-TT.